OpenStudy (anonymous):

The sum of 3 consecutive integers is 60. What is the value of the third integer?

OpenStudy (jlastino):

We can express the first number as x and the next numbers as x+1, and x+2 (since they're consecutive so x+x+1+x+2 = 60 3x+3=60 3x=60-3 3x=57 x=19 x+1=20 x+2= 21 The three numbers are 19, 20, and 21